# Nonstandard Calderón-type theorems

### v1

@@ Abstract.
This research was supported in part by the grant 23-07420S of the Czech Science Foundation, grant UNCE24/SCI/005 of Charles University, GA UK no. 148125 of Charles University and grant SVV-2025-260827
@@ 1. Introduction
Let $(\mathcal{R},\mu)$ and $(\mathcal{S},\nu)$ be non-atomic $\sigma$-finite measure spaces. The standard Marcinkiewicz interpolation theorem, first introduced by J. Marcinkiewicz in 1939 and later generalized by A. Zygmund in 1956 [Zy56], asserts that if a quasilinear operator $T$ satisfies a pair of end-point estimates,
@@
for some $\theta\in(0,1)$. Here $L^{p,q}$ stands for standard Lorentz space, see (2.2) for precise definition of its governing functional. Operators which satisfy (1.1) are said to be of weak type $(p_{0},q_{0})$ and $(p_{1},q_{1})$. This result was furthermore generalized to what is now known as Calderón’s fundamental interpolation theorem in [Ca1966]. In essence, it describes the behaviour of operators satisfying end-point estimates (1.1) on general rearrangement-invariant spaces in terms of boundedness of the so-called Calderón operator on their representation spaces. The Calderón operator is intrinsically connected to (1.1) and is defined by
@@
where $m=\frac{\frac{1}{q_{0}}-\frac{1}{q_{1}}}{\frac{1}{p_{0}}-\frac{1}{p_{1}}}$ is the slope of the corresponding interpolation segment. Using Peetre’s $K$-functional formula and Holmstedt’s formulae, one can show that for any quasilinear operator $T$ satisfying (1.1), we have $\left(Tf\right)^{*}(t)\lesssim S_{m}f^{*}(t)$ for every $t>0$, a key estimate for establishing fine boundedness properties of $T$ on various function spaces, where the star denotes the operation $f\mapsto f^{*}$ of the non-increasing rearrangement of a given function (see definitions below). It should be noted that, in the classical situation, the Calderón operator is a sum of two integral type operators involving power weights. A unified treatise of weak-type operators was conducted for example in [BeRu80].
Calderón’s interpolation theorem has since been adapted into various nonstandard settings. This was motivated by the fact that several operators of great importance such as, for instance, the fractional maximal operator or the weighted Copson-type operator governing via a certain reduction principle optimal Sobolev embeddings have (different) nonstandard endpoint behaviours. Calderón-type theorems for this sort of operators were established in [GoPi2009], in which one pair of weak-type estimates has been substituted by either boundedness from $L^{1}$ to $L^{p,1}$ or from $L^{p,\infty}$ to $L^{\infty}$. A considerably more general situation, albeit using similar approach, was considered in [Mal:12]. Completely different techniques were used in [BaGoMiPi22], where a Calderón-type theorem was established for a special integral operator that reflects interpolation properties and embeddings of Gaussian-Sobolev spaces into Orlicz-Karamata spaces.
The results of [GoPi2009] revealed a surprising phenomenon. For operators that take boundedly $L^{p_{0},1}$ to $L^{q_{0},\infty}$ and simultaneously $L^{p_{1},\infty}$ to $L^{\infty}$, the Calderón operator is once again written in the form of two suboperators, but this time one of the summands is a supremum-type operator (hence, remarkably, not linear). More precisely, the Calderón operator in these circumstances reads as
@@
It has recently been observed in [CiPiSl20201, CiPiSl20202] while investigating trace embedding theorems and Sobolev embeddings involving Frostman measures that when one of the target spaces is $L^{\infty}(\mathcal{S},\nu)$, one can, in fact, enhance the information obtained from interpolation. This improvement is captured by utilizing a structure denoted as $Y^{\langle p\rangle}(\mathcal{S},\nu)$ where $Y(\mathcal{S},\nu)$ is a rearrangement-invariant space. More precisely, these spaces are defined as the collection of all $\nu$-measurable functions $f$ satisfying
@@
Detailed treatise of these spaces has been carried out in [Tu23]. What is of importance to us is that these $Y^{\langle p\rangle}(\mathcal{R},\mu)$ spaces continuously embed into $Y(\mathcal{R},\mu)$. Moreover, these inclusions may be strict, and hence we gain an improvement of the information obtained through interpolation.
A Calderón-type interpolation theorem utilizing the structure of $Y^{\langle p\rangle}(\mathcal{R},\mu)$ spaces has recently been established in [MiPiSp]. Therein, the operators in question satisfy
@@
This research was motivated by the recent result of [Kor:24], in which borderline case of a potential operator was investigated in connection with the Luzin $N$-property and the Morse-Sard property. The respective Calderón operator for operators satisfying (1.3) is defined by
@@
The motivation for studying such operators stems from the recent research of the action of the fractional maximal operator and the Riesz potential on spaces endowed with the capacitary measures and the Hausdorff content and their important applications, see [Kor:24, MiPiSp] and the references therein.
@@ 2. Preliminaries
rearrangement-invariant Banach function spaces. For more details, see [BS].
@@
The Holmstedt formulae (see [BS, Chapter 5, Theorem 1.9 and Theorem 2.1]) assert that if $1<p_{0}<q<\infty$ and $1\leq q_{0}\leq\infty$, then for $\frac{1}{\alpha}=\frac{1}{p_{0}}-\frac{1}{q}$ we have
@@ 3. Basic functional properties and examples
The implication $(i)\Rightarrow(ii)$ is an immediate consequence of the $K$-inequality ([BS, Chapter 5, Theorem 1.11]). Let us now show the converse implication. First, by the Holmstedt formula, we have
@@
and observe that the quantity in parenthesis is non-increasing. Hence, by [GoPi2009, Lemma 3.1 (ii)] with $\beta=0$ and $\alpha=-\frac{1}{p_{1}}+\frac{r}{p_{0}}>0$ in their notation, we have
@@
Now, if $s_{2}<\infty$, utilizing (3.9) once more, [GoOpPi, Theorem 3.2 (i)] tells us that
@@ 4. Main results
Let now $g=g^{*}\in\bar{X}(0,\mu(\mathcal{R}))$ be fixed. Without loss of generality, we may assume that $\rm{supp}\ g$ is bounded. Since $(\mathcal{R},\mu)$ is nonatomic, we find $g_{0}\in\mathcal{M}^{+}(\mathcal{R},\mu)$ be such that $g^{*}=g_{0}^{*}$. By virtue of [BS, Chapter 3, Theorem 2.13] we find $T_{1}\colon(L^{1}+L^{\infty})(\mathcal{R},\mu)\to(L^{1}+L^{\infty})(0,\infty)$ satisfying $T_{1}f=\chi_{(0,\mu(\mathcal{R}))}T_{1}f$ for every $f\in(L^{1}+L^{\infty})(\mathcal{R},\mu)$, $\max\{\|T_{1}\|_{L^{1}(\mathcal{R},\mu)\to L^{1}(0,\infty)},\|T_{1}\|_{L^{\infty}(\mathcal{R},\mu)\to L^{\infty}(0,\infty)}\}\leq 1$ and $T_{1}g_{0}=g^{*}$. Furthermore, the second property, by virtue of [BS, Chapter 3, Theorem 2.2], implies that $\|T\|_{Z(\mathcal{R},\mu)\to Z(0,\infty)}\leq 1$ for every rearrangement-invariant space $Z$. We further define
@@
and $T_{2}f(t)=\tilde{T}_{2}f(t)\chi_{(0,\nu(S))}(t)$ for every $f\in\mathcal{(}L^{1}+L^{\infty})(0,\infty)$ and $t>0$. Let us observe that $(\tilde{T}_{2}\circ T_{1}f)(t)\leq t^{\frac{rp_{1}-p_{0}}{p_{0}p_{1}}}(T_{1}f)^{**}(t^{r})$ for every $f\in\mathcal{M}(\mathcal{R},\mu)$ and $t>0$. We now find a function $h\in\mathcal{M}^{+}(\mathcal{S},\nu)$ satisfying $h^{*}=(T_{2}g)^{*}$. Now, since the support of $g$ is bounded, $g\in L^{1}(0,\left|{\rm{supp}\ g}\right|)$. This implies that $h^{*}(t)\to 0,t\to\infty$. Thus, by [BS, Chapter 2, Corollary 7.6] we may find a measure-preserving transformation $\sigma\colon\rm{supp}\ h\to\rm{supp}\ h^{*}$ such that $h=(h^{*}\circ\sigma)\chi_{\{\rm{supp}\ h\}}$ $\nu$-a.e. Setting $T_{3}f_{0}(x)=f_{0}(\sigma(x))\chi_{\{\rm{supp}\ h\}}$ for $f_{0}\in\mathcal{M}(0,\infty)$ and $x\in\mathcal{S}$, we have by [BS, Chapter 2, Proposition 7.2] that $(T_{3}f_{0})^{*}=(f_{0}\chi_{\{\rm{supp}\ h\}})^{*}$ a.e. in $(0,\infty)$. Finally, we define the operator $T\colon\mathcal{M}(\mathcal{R},\mu)\to\mathcal{M}(\mathcal{S},\nu)$ by
@@
Here, the second inequality is [BS, Theorem 3.2 (i)] with weights
@@
Now, by [GoPi2009, Lemma 3.1 (ii)], we have that
@@
where the penultimate line comes from [GoPi2009, Lemma 3.1.(i)]. For the remaining case $q_{1}=\infty$, by Remark 4.4 we have $Y_{q_{0},\infty}f=R_{q_{0}}f$ for every $f\in\mathcal{M}(0,\infty)$. Thus, the result follows from Proposition 3.6. ∎
@@
Here the $\lesssim$” is [GoPi2009, Lemma 3.1. (ii)] with

### v2

@@ Abstract.
This research was supported in part by the grant 23-04720S of the Czech Science Foundation, grant UNCE24/SCI/005 of Charles University, GA UK no. 148125 of Charles University and grant SVV-2025-260827
@@ 1. Introduction
Let $(\mathcal{R},\mu)$ and $(\mathcal{S},\nu)$ be non-atomic $\sigma$-finite measure spaces. The standard Marcinkiewicz interpolation theorem, first introduced by J. Marcinkiewicz in 1939 and later generalized by A. Zygmund in 1956 [13], asserts that if a quasilinear operator $T$ satisfies a pair of end-point estimates,
@@
for some $\theta\in(0,1)$. Here $L^{p,q}$ stands for standard Lorentz space, see (2.2) for precise definition of its governing functional. Operators which satisfy (1.1) are said to be of weak type $(p_{0},q_{0})$ and $(p_{1},q_{1})$. This result was furthermore generalized to what is now known as Calderón’s fundamental interpolation theorem in [4]. In essence, it describes the behaviour of operators satisfying end-point estimates (1.1) on general rearrangement-invariant spaces in terms of boundedness of the so-called Calderón operator on their representation spaces. The Calderón operator is intrinsically connected to (1.1) and is defined by
@@
where $m=\frac{\frac{1}{q_{0}}-\frac{1}{q_{1}}}{\frac{1}{p_{0}}-\frac{1}{p_{1}}}$ is the slope of the corresponding interpolation segment. Using Peetre’s $K$-functional formula and Holmstedt’s formulae, one can show that for any quasilinear operator $T$ satisfying (1.1), we have $\left(Tf\right)^{*}(t)\lesssim S_{m}f^{*}(t)$ for every $t>0$, a key estimate for establishing fine boundedness properties of $T$ on various function spaces, where the star denotes the operation $f\mapsto f^{*}$ of the non-increasing rearrangement of a given function (see definitions below). It should be noted that, in the classical situation, the Calderón operator is a sum of two integral type operators involving power weights. A unified treatise of weak-type operators was conducted for example in [2].
Calderón’s interpolation theorem has since been adapted into various nonstandard settings. This was motivated by the fact that several operators of great importance such as, for instance, the fractional maximal operator or the weighted Copson-type operator governing via a certain reduction principle optimal Sobolev embeddings have (different) nonstandard endpoint behaviours. Calderón-type theorems for this sort of operators were established in [8], in which one pair of weak-type estimates has been substituted by either boundedness from $L^{1}$ to $L^{p,1}$ or from $L^{p,\infty}$ to $L^{\infty}$. A considerably more general situation, albeit using similar approach, was considered in [10]. Completely different techniques were used in [1], where a Calderón-type theorem was established for a special integral operator that reflects interpolation properties and embeddings of Gaussian-Sobolev spaces into Orlicz-Karamata spaces.
The results of [8] revealed a surprising phenomenon. For operators that take boundedly $L^{p_{0},1}$ to $L^{q_{0},\infty}$ and simultaneously $L^{p_{1},\infty}$ to $L^{\infty}$, the Calderón operator is once again written in the form of two suboperators, but this time one of the summands is a supremum-type operator (hence, remarkably, not linear). More precisely, the Calderón operator in these circumstances reads as
@@
It has recently been observed in [5, 6] while investigating trace embedding theorems and Sobolev embeddings involving Frostman measures that when one of the target spaces is $L^{\infty}(\mathcal{S},\nu)$, one can, in fact, enhance the information obtained from interpolation. This improvement is captured by utilizing a structure denoted as $Y^{\langle p\rangle}(\mathcal{S},\nu)$ where $Y(\mathcal{S},\nu)$ is a rearrangement-invariant space. More precisely, these spaces are defined as the collection of all $\nu$-measurable functions $f$ satisfying
@@
Detailed treatise of these spaces has been carried out in [12]. What is of importance to us is that these $Y^{\langle p\rangle}(\mathcal{R},\mu)$ spaces continuously embed into $Y(\mathcal{R},\mu)$. Moreover, these inclusions may be strict, and hence we gain an improvement of the information obtained through interpolation.
A Calderón-type interpolation theorem utilizing the structure of $Y^{\langle p\rangle}(\mathcal{R},\mu)$ spaces has recently been established in [11]. Therein, the operators in question satisfy
@@
This research was motivated by the recent result of [9], in which borderline case of a potential operator was investigated in connection with the Luzin $N$-property and the Morse-Sard property. The respective Calderón operator for operators satisfying (1.3) is defined by
@@
The motivation for studying such operators stems from the recent research of the action of the fractional maximal operator and the Riesz potential on spaces endowed with the capacitary measures and the Hausdorff content and their important applications, see [9, 11] and the references therein.
@@ 2. Preliminaries
rearrangement-invariant Banach function spaces. For more details, see [3].
@@
The Holmstedt formulae (see [3, Chapter 5, Theorem 1.9 and Theorem 2.1]) assert that if $1<p_{0}<q<\infty$ and $1\leq q_{0}\leq\infty$, then for $\frac{1}{\alpha}=\frac{1}{p_{0}}-\frac{1}{q}$ we have
@@ 3. Basic functional properties and examples
The implication $(i)\Rightarrow(ii)$ is an immediate consequence of the $K$-inequality ([3, Chapter 5, Theorem 1.11]). Let us now show the converse implication. First, by the Holmstedt formula, we have
@@
and observe that the quantity in parenthesis is non-increasing. Hence, by [8, Lemma 3.1 (ii)] with $\beta=0$ and $\alpha=-\frac{1}{p_{1}}+\frac{r}{p_{0}}>0$ in their notation, we have
@@
Now, if $s_{2}<\infty$, utilizing (3.9) once more, [7, Theorem 3.2 (i)] tells us that
@@ 4. Main results
Let now $g=g^{*}\in\bar{X}(0,\mu(\mathcal{R}))$ be fixed. Without loss of generality, we may assume that $\rm{supp}\ g$ is bounded. Since $(\mathcal{R},\mu)$ is nonatomic, we find $g_{0}\in\mathcal{M}^{+}(\mathcal{R},\mu)$ be such that $g^{*}=g_{0}^{*}$. By virtue of [3, Chapter 3, Theorem 2.13] we find $T_{1}\colon(L^{1}+L^{\infty})(\mathcal{R},\mu)\to(L^{1}+L^{\infty})(0,\infty)$ satisfying $T_{1}f=\chi_{(0,\mu(\mathcal{R}))}T_{1}f$ for every $f\in(L^{1}+L^{\infty})(\mathcal{R},\mu)$, $\max\{\|T_{1}\|_{L^{1}(\mathcal{R},\mu)\to L^{1}(0,\infty)},\|T_{1}\|_{L^{\infty}(\mathcal{R},\mu)\to L^{\infty}(0,\infty)}\}\leq 1$ and $T_{1}g_{0}=g^{*}$. Furthermore, the second property, by virtue of [3, Chapter 3, Theorem 2.2], implies that $\|T\|_{Z(\mathcal{R},\mu)\to Z(0,\infty)}\leq 1$ for every rearrangement-invariant space $Z$. We further define
@@
and $T_{2}f(t)=\tilde{T}_{2}f(t)\chi_{(0,\nu(S))}(t)$ for every $f\in\mathcal{(}L^{1}+L^{\infty})(0,\infty)$ and $t>0$. Let us observe that $(\tilde{T}_{2}\circ T_{1}f)(t)\leq t^{\frac{rp_{1}-p_{0}}{p_{0}p_{1}}}(T_{1}f)^{**}(t^{r})$ for every $f\in\mathcal{M}(\mathcal{R},\mu)$ and $t>0$. We now find a function $h\in\mathcal{M}^{+}(\mathcal{S},\nu)$ satisfying $h^{*}=(T_{2}g)^{*}$. Now, since the support of $g$ is bounded, $g\in L^{1}(0,\left|{\rm{supp}\ g}\right|)$. This implies that $h^{*}(t)\to 0,t\to\infty$. Thus, by [3, Chapter 2, Corollary 7.6] we may find a measure-preserving transformation $\sigma\colon\rm{supp}\ h\to\rm{supp}\ h^{*}$ such that $h=(h^{*}\circ\sigma)\chi_{\{\rm{supp}\ h\}}$ $\nu$-a.e. Setting $T_{3}f_{0}(x)=f_{0}(\sigma(x))\chi_{\{\rm{supp}\ h\}}$ for $f_{0}\in\mathcal{M}(0,\infty)$ and $x\in\mathcal{S}$, we have by [3, Chapter 2, Proposition 7.2] that $(T_{3}f_{0})^{*}=(f_{0}\chi_{\{\rm{supp}\ h\}})^{*}$ a.e. in $(0,\infty)$. Finally, we define the operator $T\colon\mathcal{M}(\mathcal{R},\mu)\to\mathcal{M}(\mathcal{S},\nu)$ by
@@
Here, the second inequality is [3, Theorem 3.2 (i)] with weights
@@
Now, by [8, Lemma 3.1 (ii)], we have that
@@
where the penultimate line comes from [8, Lemma 3.1.(i)]. For the remaining case $q_{1}=\infty$, by Remark 4.4 we have $Y_{q_{0},\infty}f=R_{q_{0}}f$ for every $f\in\mathcal{M}(0,\infty)$. Thus, the result follows from Proposition 3.6. ∎
@@
Here the $\lesssim$” is [8, Lemma 3.1. (ii)] with